OpenStudy (anonymous):

i will challenge you to answer this question... 4int_{0}^{1}int_{0}^{sqrt{1-x}}int_{4^{2}+2y ^{2}}^{2-x ^{2}}dzdydx

OpenStudy (amistre64):

helps i fyou code it right ...

OpenStudy (amistre64):

accursed keyboards lol

OpenStudy (anonymous):

the answer is 180 degrees who can prove

OpenStudy (apoorvk):

haha.. please 'draw' the question

OpenStudy (amistre64):

there is no theta in it; how can it be answered by degees?

OpenStudy (apoorvk):

okay he means x. and i think its an integration.. bert plz repost the question am sure someone can do it.

OpenStudy (amistre64):

me thinks you have a problem, looked up an answer; and want us to fill in the details for you .... thats my suspicions

OpenStudy (amistre64):

but to even begin, you need to give us a better read on what it is your trying to integrate

OpenStudy (anonymous):

use the trigonometric identities, transformation by parts, integration by parts... all that is usable

OpenStudy (anonymous):

ok i will going rewrite it

OpenStudy (anonymous):

\[4\int\limits_{0}^{1}\int\limits_{0}^{\sqrt{1-x ^{2}}}\int\limits_{x^{2}+2y}^{2-x ^{2}}\]

OpenStudy (anonymous):

dzdydx